# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 6.1 Various models
Besides finite size strangelets and strange stars, also a branch of strange white dwarfs was also discussed Glendenning et al. (1995). Such dwarfs would have conventional nuclear crusts and strange quark matter cores. Possible candidates for strange dwarfs were suggested in Ref. Kurban et al. (2022). Potential implications for gravitational wave astronomy are highlighted in Ref. Perot et al. (2023).
@@
Recently, a novel state of deconfined matter was proposed in Ref. Zhitnitsky and Maroudas (2025) using the model described in Ref. Zhitnitsky (2003). This proposed object consists of a core of quarks (antiquarks) in a color-superconducting state surrounded by an electrosphere of positrons (electrons), and enclosed by an axion domain wall. Large axion domain walls can form large bubbles filled with $u$, $d$ and $s$ quarks. Earlier, the idea that soliton-like configurations could serve as dark matter arose in Refs. Friedberg et al. (1976); Coleman (1985); Lee (1987); Lee and Pang (1989, 1992). Among other conditions, the possibility of existence of these objects is based on the feasible existence of a conserved charge associated with an ungauged unbroken continuous internal symmetry. In Ref. Coleman (1985) such objects were called $Q$ balls. Large $Q$ balls may form so-called boson stars. A review of these topics can be found in Shnir (2018); Nugaev and Shkerin (2020). Dark matter can be associated with a light boson that has a mass $m\sim 10-22$ eV and a de Broglie wavelength of $\sim 1$ kpc. This is often called the fuzzy dark matter. A soliton is surrounded by an envelope resembling a cold dark matter hallo Hui et al. (2017).
@@
We should mention that recently, the mass and radius of the Type-I X-ray burster and accretion-powered millisecond pulsar XTE J1814-138 were inferred to be $M=(1.21\pm 0.05)M_{\odot}$ and $R=7.0\pm 0.4$ km (68.3% credibility interval), using the pulse profile modeling technique Kini et al. (2024). If were confirmed, the compact object of such a small radius cannot be described by conventional purely hadronic equations of state of the neutron stars. This could then be a candidate for the strange star, the pion condensate nucleus-star, the Lee-Wick $\sigma$ condensate compact object, or something else.
6.2 Cooling of strange stars and strangelets
@@
The neutrino emissivity of quark matter was first considered by Iwamoto in Ref. Iwamoto (1982). In Ref. Pizzochero (1991) the surface temperature of the strange star was found to be lower than that of a neutron star of the same age (ordinary or with mesonic condensates). The effect of color superconductivity on the cooling of strange and hybrid stars has been studied in Refs. Blaschke et al. (2000, 2001); Grigorian et al. (2005) and subsequent works. Some models show that cooling can be so fast that strange stars may presently have temperatures smaller than the cosmic microwave background radiation Shaulov et al. (2022); Kupriyanova and Shaulov (2024).
@@
Since we have mentioned the axion-based model of nuggets, we should also mention a possible axion contribution to cooling of neutron stars. The emission of axions by neutron stars was, first, discussed in Ref. Iwamoto (1984) without the inclusion of $NN$ interactions. Then in Ref. Voskresensky and Senatorov (1986) it was shown that axion emission should be strongly affected by in-medium baryon-baryon interactions similarly to the corresponding reaction of the neutrino production.
7 Dilute stable nuclear systems
7.1 Again scalar condensate
@@
The analysis of two independent nuclear emulsion experiments exposed to beams of 16O and 56Fe ions with 2A GeV kinetic energy allowed for the interpretation of some events as the relatively rare occurrence of anomalous nuclear fragments interacting with an unexpectedly large cross section as described in Ref. Friedlander et al. (1983). These objects were called anomalons Bayman and Tang (1987). Although these results were not reproduced by subsequent measurements, uncertainty remains.
@@
$G_{N}(\epsilon,\vec{p},n)=\frac{a}{\epsilon-\xi_{\vec{p}}+i\,0\,{\rm sign}\epsilon}\,,\quad\xi_{\vec{p}}=\frac{p^{2}-p_{\rm F}^{2}}{2\,m_{N}^{*}}\approx v_{\rm F}(p-p_{\rm F})\,,\quad v_{\rm F}=p_{\rm F}/m_{N}^{*}\,.$ (31)
@@
where $E_{N}$ is the energy density of the nucleon system. Parameter $K_{N}$ is related to the standard compressibility modulus of the nuclear matter at saturation that we mentioned in Section 2 as $K=9K_{N}$.
The particle-hole scattering amplitude in the spin zero channel on the Fermi surface can be written throught a propagator of the scalar bosonic mode Kolomeitsev and Voskresensky (2016):
@@
$\intop_{-\infty}^{+\infty}\frac{{\rm d}\epsilon}{2\pi i}\intop\frac{2{\rm d}^{3}p}{(2\pi)^{3}}G_{+}G_{-}=-a^{2}N\Phi(s,x)\,,\,\,s=\frac{\omega}{kv_{\rm F}}\,,\,\,x=\frac{k}{p_{\rm F}}.$ (34)
@@
which is valid for $1-|f_{0}|\ll 1$ and $k\ll p_{\rm F}$. If $f_{0}<-1$, i.e. the compressibility of the system becomes negative according to Eq. (32), one has $\Im\omega_{\rm d}(k)>0$, and the scalar excitation modes grow exponentially with time as $\propto\exp(+\Im\omega_{\rm d}t)$. This instability is called the Pomeranchuk instability of a Fermi liquid. It may result in the spinodal instability which appears in systems with a van der Waals-like equations of state. Another possibility is the formation of the Bose condensate of a scalar field, in analogy with the Lee-Wick state, but in dilute matter Kolomeitsev and Voskresensky (2016).
We assume that growing unstable bosonic modes may form the Bose condensate described by a scalar boson complex field $\phi$. Such a scalar field can be introduced by means of the Hubbard–Stratonovich transformation or by a formal replacement of the contact interaction with an exchange of an artificial heavy scalar boson Kolomeitsev and Voskresensky (2016). The effective Lagrangian density of the condensate field, taken in the simplest form $\phi=\phi_{0}e^{-i\omega_{c}t+i\vec{k}_{0}\vec{r}}$, can be presented as
@@
It was evaluated in Brovman and Kagan (1972); Brovman and Kholas (1974) and Dyugaev (1982) in case of one-component fermion system. Energetically preferable is the state with $\omega_{c}=0$ as argued in Ref. Kolomeitsev and Voskresensky (2016). Then we can estimate for $k_{0}\ll p_{\rm F}$,
@@
$a^{2}N\widetilde{\omega}^{2}_{s}(k_{0})\phi_{0}+\Lambda(0,k_{0})\,|\phi_{0}|^{2}\phi_{0}=0\,,\quad\widetilde{\omega}^{2}_{s}(k_{0})=|f_{0}(k_{0})|^{-1}-\Re\Phi(0,k_{0})\,.$ (42)
@@
$|\phi_{0}|^{2}=-\frac{N\widetilde{\omega}^{2}_{s}(k_{0})}{a^{2}\lambda}\,\Big(1+\frac{k_{0}^{2}}{2p_{\rm F}^{2}}\Big)^{-1}\,,\quad E_{\rm B}=-\frac{N^{2}\widetilde{\omega}^{4}_{s}(k_{0})}{2\lambda}\Big(1+\frac{k_{0}^{2}}{2p_{\rm F}^{2}}\Big)^{-1}.$ (43)
@@
Reference Kolomeitsev and Voskresensky (2016) employed a simple phenomenological parametrization for the volume part of the energy per nucleon, $\mathcal{E}_{N\rm}$ as an expansion in powers of $p_{\rm F}/m_{N}$ proposed in Ref. Kaiser et al. (2002),

### v2

@@ VI.1 Quark models
Besides finite size strangelets and strange stars, also a branch of strange white dwarfs was also discussed [104]. Such dwarfs would have conventional nuclear crusts and strange quark matter cores. Possible candidates for strange dwarfs were suggested in Ref. [169]. Potential implications for gravitational wave astronomy are highlighted in Ref. [221].
@@
Recently, other state of deconfined matter was proposed in Ref. [301] based on the model described in Ref. [302]. The compact soliton-like object consists of a core of quarks (antiquarks) in a color-superconducting state surrounded by an electrosphere of positrons (electrons), and enclosed by an axion domain wall. Large axion domain walls can form large bubbles filled with $u$, $d$ and $s$ quarks. Earlier, the idea that soliton-like configurations could serve as dark matter arose in Refs. [90, 55, 184, 178, 177]. Among other conditions, the possibility of existence of these objects is based on the feasible existence of a conserved charge associated with an ungauged unbroken continuous internal symmetry. In Ref. [55] such objects were called $Q$ balls. Large $Q$ balls may form so-called boson stars. There is a scenario that the dark matter is associated with slow light bosons with a mass $m\sim 10-22$ eV and the huge de Broglie wavelength of $\sim 1$ kpc. This is called the fuzzy dark matter. A soliton is surrounded by an envelope resembling a cold dark matter hallo [126]. Recent work [201] proposes a model for the cosmological formation of superheavy $Q$ balls as MACHOs (massive compact halo objects), the objects in the mass range $(10^{-7}-10^{-6})\,M_{\odot}$. Also it shows that it is possible to produce asteroid mass $Q$ balls. A review of these topics can be found in [262, 219, 201].
Finally, if the mass distribution of strangelets is obtained, it is possible to formulate well-focused campaigns to search for these objects, cf. [67].
The most promising signals of the scalar and pion condensate nuclei-stars and strange stars seem to be an abnormal mass-radius relation [210, 37], an $r$ mode spin-down [191, 71, 158], and specific stellar cooling properties [292, 132, 227, 32, 30, 111].
VI.2 Abnormal $\pi$ condensate and strange compact stars and the star mass-radius relation
For quite some time there has been a consensus that most neutron stars being produced in supernova explosions have masses near $1.4M_{\odot}$ close to the neutron star maximum mass $M_{\rm max}\simeq 1.5M_{\odot}$ [268, 43]. For $M>M_{\rm max}$ one deals with the black hole. Most of the existing that time hadronic equations of state could describe neutron stars with such masses. However then compact stars of different masses were observed. By now it is well established that the compact star masses vary in a broad interval [174]. At present the minimal value of the accurately measured mass is $1.174(4)\,M_{\odot}$ for the pulsar PSR J0453+1559 [194]. There was the claim that the object HESS CCO XMMU J173203.3-344518 has a very small mass of $0.77^{+0.20}_{-0.17}M_{\odot}$ [70]. The analysis of Ref. [70] was criticized in Ref. [8] with the conclusion that a small mass source is not needed for the description of observed spectra. Thereby one should consider the result  [70] with a caution.
@@ VI.2 Abnormal $\pi$ condensate and strange compact stars and the star mass-radius relation
Today, the highest precisely-measured mass of a neutron star is $2.08(7)\,M_{\odot}$ for PSR J0740+6620 [57, 87]. Recent analyses of the NICER data indicate that there is no significant variation of star radii for stars with masses varying between 1.4 and 2.0 $M_{\odot}$, $R\simeq(12-13)$ km. A more detailed discussion can be found in Ref. [161]. Although any strong softening of the equation of state leads to a decrease of the radius for the objects with fixed mass, e.g. see Ref. [129], the maximum mass of a pion-condensate nucleus-star, a hybrid star, or a strange star can be smaller or larger than the purely hadronic star, depending on the utilized model. In this respect we should mention that recently, the mass and radius of the Type-I X-ray burster and accretion-powered millisecond pulsar XTE J1814-138 were inferred in Ref. [145] to be $M=(1.21\pm 0.05)\,M_{\odot}$ and $R=7.0\pm 0.4$ km ( with 68.3% credibility). If confirmed, the compact object of such a small radius could not be described by a conventional purely hadronic equation of state. Then, this object would be a candidate for a strange star, a pion condensate nucleus-star, a compact object with the Lee-Wick $\sigma$ condensate, or something else.
Reference [207] suggested that a massive cold neutron star can undergo the first-order phase transition to the pion condensate state with a blowing off a part of matter or at least a giant glitch. Then a similar idea was applied to the first-order phase transition to a hybrid (quark-hadron) star configuration [139]. In majority of papers the later term “hybrid star” is associated with the star heaving a quark interior and an extended hadron exterior although consequences of any strong first-order phase transition are approximately the same, see further discussion in [210]. Then, an interesting idea appeared that a neutron star and a star, where the first-order phase transition to the state with a superdense interior (pion condensate, $\sigma$ condensate, quark matter or something else) occurred, can have the same mass at very different radii. If mass-twins were observed, it would be an evidence for the existence of compact stars different from ordinary neutron stars (e.g. hybrid stars, strange stars, pion-condensate nuclei-stars, or something else), see Refs. [99, 105, 251, 25, 197] and references therein.
VI.3 Cooling of $\pi$ condensate abnormal nuclei, nuclei-stars, strangelets and strange stars
Below we show that nuclei-stars held together by a pion condensate and neutron stars containing a broad pion condensation region in their interiors as well as hybrid stars with an extended region of a quark 2SC phase are probably rather massive rapidly cooling stars, with $M\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}1.5\,M_{\odot}$.
At present time there exist data on the surface temperatures of some pulsars as function of their age, see [229, 228]. Therefore, it is worthwhile to check whether existing models can describe them. The data can be separated on three groups: slow, intermediate, and rapid coolers. The data impose that the averaged neutrino emissivity for the slow and rapid coolers should differ by factor $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{3}$. The minimal cooling scenario, in which important in-medium effects are disregarded, cannot appropriately describe the data. Shortcomings of this scenario are discussed in Refs. [292, 31, 108].
The most efficient processes of the neutrino production in neutron stars are one-nucleon processes because of their large phase-space volume. These are the DU processes, e.g. $n\to p+e+\bar{\nu}_{e}$ and $p+e\to n+\nu_{e}$, and the processes going on the mesonic condensates, e.g. on the pion condensate, such as $n+\pi^{-}_{\rm cond}\to n+e+\bar{\nu}_{e}$. The emissivity of the DU process can be roughly estimated as
$\epsilon_{\nu}^{\rm{DU}}\sim 10^{27}(n_{e}/n_{0})^{1/3}e^{-\Delta/T}T_{9}^{6}\Theta(n-n_{c}^{\rm{DU}})\frac{\mbox{erg}}{\mbox{cm}^{3}\cdot\mbox{s}},$ (35)
where $T_{9}$ is the temperature measured in $10^{9}$ K, $n_{e}$ is the density of electrons. The suppression factor $e^{-\Delta/T}$ appears for temperatures below the critical temperature of a nucleon pairing $T_{\Delta}$, and a pairing gap $\Delta\propto(T_{\Delta}-T)\Theta(T_{\Delta}-T)$, where $\Theta(x)$ is the step-function. The DU process occurs for $n>n_{c}^{\rm DU}$, when the fraction of protons exceeds $\sim 10\%$. The DU process is so efficient that, if it occurred, the stars with $M>M_{\rm DU}+0.1M_{\odot}$ would cool so rapidly that they would not be manifested in the cooling data [154, 147]. Therefore, the deficiency of the “minimal cooling + DU” scenario is that it would require all stars in the cooling diagram (temperature vs. time) to have a mass close to the threshold for the DU processes to occur in the star center, with an uncertainty of only $0.1M_{\odot}$. However now it is known that masses of neutron stars proved to be essentially different. Moreover, there are equations of state in the literature, in which the DU process either does not appear in the whole available density interval or it occurs only for $M>1.5M_{\odot}$. For example, in case of the famous variational equation of state [5] the threshold mass for the DU process is $M_{\rm DU}\simeq 2M_{\odot}$.
In works [247, 292, 33, 110, 29, 31, 108, 109] the “nuclear medium cooling scenario” was developed. It relied on the results of Refs. [277, 278, 279, 292] where it was assumed that neutron stars with measured surface temperatures (first data fixed only upper limits on surface temperatures) may have very different masses, and that neutrino emissivity of the star strongly depends on the nucleon density and, consequently, on the neutron star mass. This dependence appeared since the in-medium modification of the pion exchange was taken into account that significantly affected two-nucleon reaction rates. Within the “nuclear medium cooling scenario” the DU reactions were not allowed, at least for stars with masses $M<1.5M_{\odot}$.
Another process of the one-nucleon origin is a reaction going on the pion condensate, e.g. $n+\pi_{\rm cond}^{-}\to n+e+\bar{\nu}$. Its emissivity is roughly estimated as
$\epsilon_{\nu}^{\pi}\sim 10^{27}\frac{\phi^{2}}{m_{\pi}^{2}}\Big(\frac{n_{e}}{n_{0}}\Big)^{1/3}e^{-\Delta/T}T_{9}^{6}\Theta(n-n_{c}^{\pi})\frac{\mbox{erg}}{\mbox{cm}^{3}\cdot\mbox{s}},$ (36)
where $n_{c}^{\pi}\sim(1.5-4)n_{0}$ depending on the model, and $\phi$ is the amplitude of the pion condensate field, $\phi\sim(0.1-1)m_{\pi}$. This process would be too efficient as the DU process, if $n_{c}^{\pi}$ were smaller $2n_{0}$, $M\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}(1.3-1.35)M_{\odot}$. It would affect most of the neutron star population and lead to a contradiction with the measured Log(N) - Log(S) distribution. We should also mention the DU-like processes occurring on hyperons and $\Delta$ isobars at densities exceeding the corresponding critical densities. Values of these emissivities are similar to those for the $\pi$ condensate processes, cf. [112, 109]
In the presence of nucleon pairing (for $T<T_{\Delta}$) a new important process of the quasi-one-nucleon origin occurs. It is associated with breaking and formation of nucleon pairs, so the name — the pair-breaking-formation (PBF) process [85, 279, 185, 155], e.g. $n_{\rm pair}\to n+\nu+\bar{\nu}$. Its emissivity is estimated as
$\epsilon_{\nu}^{\rm{PBF}}\sim 10^{28}\big(\Delta/{\mbox{MeV}}\big)^{7}(T/\Delta)^{1/2}e^{-2\Delta/T}\frac{\mbox{erg}}{\mbox{cm}^{3}\cdot\mbox{s}}\,.$ (37)
The in-medium renormalization of the vertices is crucially important for the PBF processes [155].
Note that processes going on $\pi$ condensate and the PBF processes accounting the renormalization of the vertices are purely in-medium processes, which would not be consistent to include in the minimal cooling scheme disregarding effects of the polarization of the medium. However the PBF process was included, whereas the $\pi^{-}$ condensate processes were disregarded, even though the latter would occur already in the stars with $M\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}1M_{\odot}$, since the electron chemical potential would then reaches the bare pion mass if in-medium effects were switched off. The pionization of the neutron star matter can only be precluded if the effects of the medium on pions are taken into account [215].
In the absence of the DU process and PBF process in the minimal cooling scheme the main role in the neutron star cooling is taken by the modified Urca processes, e.g. $n+n\to n+p+e+\bar{\nu}_{e}$. Its emissivity is roughly estimated as
$\epsilon_{\nu}^{\rm{MU}}\sim 10^{21}\Big(\frac{n_{e}}{n_{0}}\Big)^{1/3}e^{-2\Delta/T}T_{9}^{8}\frac{\mbox{erg}}{\mbox{cm}^{3}\cdot\mbox{s}}.$ (38)
Within the “nuclear medium cooling scenario” it was shown that the emissivity of two-nucleon reactions, e.g. $n+n\to n+p+e+\bar{\nu}_{e}$, — the medium-modified Urca processes (MMU) — significantly increases with an increasing of the central density (and the mass) of the object due to strong modification of the $NN$ interaction as demonstrated in Fig. 26. Already in Refs. [277, 278, 279] it was suggested that (i) the central densities (masses) of cooling neutron stars are different, and (ii) the nucleon-nucleon scattering is significantly enhanced in denser matter, cf. Eq. (31). The appropriate formalism for calculation of the reaction rates in medium is the non-equilibrium closed diagram formalism employing the exact Green’s functions. It was developed in the application to the neutrino emmisivity problem in Ref. [279] using the quasiparticle approximation and then in Ref. [149] was extended by inclusion of finite nucleon widths, see also [292]. In the absence of the DU processes, the MMU processes are dominant already for $n>n_{c}^{(1)}$, where $n_{c}^{(1)}<n_{0}$, see Fig. 6. Its emissivity is roughly estimated as
$\epsilon_{\nu}^{\rm{MMU}}\sim 10^{21}\frac{F(n)}{F(n_{0})}\Big(\frac{n_{e}}{n_{0}}\Big)^{1/3}e^{-2\Delta/T}T_{9}^{8}\frac{\mbox{erg}}{\mbox{cm}^{3}\cdot\mbox{s}},$ (39)
where the factor $F(n)$ appears due to a strong density dependence of the $NN$ interaction amplitude for $n>n_{c}^{(1)}$,
$F(n)\simeq 3\Big(\frac{n}{n_{0}}\Big)^{10/3}\frac{[\Gamma(n)/\Gamma(n_{0})]^{6}}{[\tilde{\omega}(k_{0}(n))/m_{\pi}]^{8}}.$ (40)
Here $\Gamma^{-1}\simeq 1+C(n/n_{0})^{1/3}\Phi(0,k_{0})$ is the $NN$ correlation factor, see (27), $\Phi(0,k_{0})\simeq 1$, and $C\simeq 1.4-1.6$ is expressed via the Landau-Migdal parameter in the spin channel, $\tilde{\omega}(k_{0}(n))$ is the effective pion gap introduced in (30). Since $\tilde{\omega}(k_{0}(n))$ decreases with increasing density one easily recovers the required factor from one to $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{3}$ in emissivity for stars with masses from 1 $M_{\odot}$ to 2 $M_{\odot}$ required to describe the cooling data, appropriately. The pion condensation is not needed for that but, if occurred at $n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(2.5-3)n_{0}$, it would not contradict the cooling data. The neutron stars with the pion condensate in their interiors and the pion condensate nuclei-stars are then rather massive rapidly cooling stars with $M\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}1.5M_{\odot}$.
A modification of the reaction rates due to multiple nucleon scattering, — the Landau-Pomeranchuk-Migdal effect,— was evaluated in Ref. [149]. Accounting for it, the neutrino emissivities acquire extra pre-factors of the type
$\frac{\omega_{\nu}^{2}}{\omega_{\nu}^{2}+\Gamma_{N}^{2}(n,T)}\,$, (41)
where $\omega_{\nu}$ is the neutrino energy and $\Gamma_{N}$ is the nucleon width, $\Gamma_{N}\sim\epsilon_{\rm F}(1-\pi^{2}T^{2}/\epsilon_{\rm F}^{2})$ and $\epsilon_{\rm F}$ is the nucleon Fermi energy. The pre-factor essentially deviates from unity only for the neutrino energy $\omega_{\nu}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}\Gamma_{N}$, with $\omega_{\nu}\sim(3-5)T$ depending on the reaction type. Thus, the multiple scattering effects are important for the description of mergers and the initial stage of the cooling of the compact stars [149, 292, 156, 10].
There are two stages in the cooling of compact objects: the initial stage, when neutrinos are trapped, and the stage of a long-time cooling when neutrinos are radiated from the compact object by the direct reactions, cf. [210]. At the initial stage of the compact star evolution (lasts from minutes to several hours depending on the mass of the object and the model) the neutrinos and antineutrinos are captured inside the star until their mean-free path will not reach the radius of the object. In case of hybrid stars there are two quantities playing the role of $R$. One is the radius of the quark or pion or sigma condensate interior core and other is the radius of the object (the core plus the hadronic exterior). The mean free paths of the neutrino and antineutrino are determined by the most efficient processes. Roughly in the process under consideration the ratios of the inverse mean free pathes are proportional to the ratio of the emissivities [242]. As function of the temperature the emissivity of the DU process is $\propto T^{6}$, see Eq. (35) and for the MU and MMU processes it is $\propto T^{8}$, see Eq. (39) and (40). The ratio of the emissivities of the DU and MU processes is $\epsilon_{\rm DU}/\epsilon_{\rm MU}\sim 10^{5}/T_{9}^{2}$. The emissivity of the MU process exceeds that of the DU one for $T\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}30$ MeV. For the in-medium modified Urca processes, MMU, the emissivity increases sharply with grows of the density. Therefore the MMU processes become more efficient than the DU processes at a smaller temperature. For the compact stars with $T>T_{\rm opac}(R)$ neutrinos and antineutrinos are trapped inside the compact star. The heat flows to the surface due to the heat conductivity. The path length of neutrinos is much larger than for quarks and hadrons and the heat conductivity is therefore governed by neutrinos. In case of hybrid stars the neutrino transport changes on the boundary between the quark and hadron phases. For the hadron phases this stage was studied in Refs. [242, 243, 93, 245, 278, 210]. In-medium effects proved to be very important, see Refs. [277, 278, 279, 210, 149, 292, 156]. As it was mentioned, appropriate optical theorem technique for the calculation the reaction rates within the nonequilibrium diagram formalism was developed within the quasiparticle approximation in [279] and with inclusion of the baryon width effects in Ref. [149]. Similar effects should be included in the description of a quark phase. With inclusion of in-medium effects associated with a decrease of the effective pion gap at an increasing baryon density, the value $T_{\rm opac}$ diminishes. For more massive neutron and hybrid stars one may expect that it is below 1 MeV.
@@ VI.3 Cooling of $\pi$ condensate abnormal nuclei, nuclei-stars, strangelets and strange stars
The neutrino emissivity of the quark matter was first calculated by Naoki Iwamoto in Ref. [132]. In Ref. [227] the surface temperature of the strange star was found to be lower than that of a neutron star of the same age (ordinary or with mesonic condensates). The effect of the color superconductivity on the cooling of strange and hybrid stars for $T<T_{opac}$ has been first studied in Refs. [32, 30, 111] and in some subsequent works. Some models show that cooling can be so fast that strange stars may presently have temperatures smaller than the cosmic microwave background radiation [261, 168].
In the case of the 2SC superconducting phase, the most efficient process is the DU process going on unpaired quarks. The cooling evolution is then very rapid that does not agree with the cooling data. Thereby, it is important to have a model, which does not allow for unpaired quarks. Since there exist other attractive quark pairing channels with a rather weak attraction ($X$ channels), Ref. [111] introduced a hypothetical 2SC+$X$ phase of two-flavor quark matter, where all the quarks are paired, some strongly in the 2SC channel and some weakly in the $X$ channel. Typical values of pairing gaps $\Delta_{X}$ in the $X$ channel should be in the range 10 keV–-1 MeV, as for example in the case of the CSL pairing [248, 252]. Nevertheless, Ref. [111] did not identify the $X$ channel with a spin one pairing channel like the CSL one simply because this phase does not coexist with the 2SC phase, which was assumed as the dominant pairing structure. Furthermore, Ref. [111] supposed that there are no Goldstone bosons in the 2SC+$X$ phase. As the result it was demonstrated that the cooling data can be appropriately described at such assumptions.
@@
Since we have mentioned the axion-based model of nuggets, we should also mention a possible axion contribution to cooling of neutron stars. The emission of axions by neutron stars was, first, discussed in Ref. [131] without the inclusion of $NN$ interactions in the hadron phase. Then in Ref. [278] it was shown that axion emission should be strongly affected by in-medium baryon-baryon interactions in the hadron phase similarly to the corresponding reaction of the neutrino production. Recent discussion of the axion contribution to the cooling of the quark matter can be found in [258].
VII Dilute stable nuclear systems
VII.1 Again, a scalar condensate
@@
The analysis of two independent nuclear emulsion experiments exposed to beams of 16O and 56Fe ions with 2A GeV kinetic energy allowed for the interpretation of some events as the relatively rare occurrence of anomalous nuclear fragments interacting with an unexpectedly large cross section as described in Ref. [91]. These objects were called anomalons [24]. Although these results were not reproduced by subsequent measurements, uncertainty remains.
@@
$G_{N}(\epsilon,\vec{p},n)$ $=\frac{a}{\epsilon-\xi_{\vec{p}}+i\,0\,{\rm sign}\epsilon}\,$$\quad$
$\xi_{\vec{p}}=\frac{p^{2}-p_{\rm F}^{2}}{2\,m_{N}^{*}}$ $\approx v_{\rm F}(p-p_{\rm F})\,,\quad v_{\rm F}=p_{\rm F}/m_{N}^{*}\,.$ (42)
@@
where $E_{N}$ is the energy density of the nucleon system. Parameter $K_{N}$ is related to the standard compressibility modulus of the nuclear matter at saturation that we mentioned in Section II as $K=9K_{N}$.
The particle-hole scattering amplitude in the spin zero channel on the Fermi surface can be written throught a propagator of the scalar bosonic mode [158]:
@@
$\intop_{-\infty}^{+\infty}\frac{{\rm d}\epsilon}{2\pi i}\intop\frac{2{\rm d}^{3}p}{(2\pi)^{3}}G_{+}G_{-}=-a^{2}N\Phi(s,x)\,$,
$\qquad s=\frac{\omega}{kv_{\rm F}}\,,\,\,x=\frac{k}{p_{\rm F}}.$ (45)
@@
which is valid for $1-|f_{0}|\ll 1$ and $k\ll p_{\rm F}$. If $f_{0}<-1$, i.e. the compressibility of the system becomes negative according to Eq. (43), one has $\Im\omega_{\rm d}(k)>0$, and the scalar excitation modes grow exponentially with time as $\propto\exp(+\Im\omega_{\rm d}t)$. This instability is called the Pomeranchuk instability of a Fermi liquid. It may result in the spinodal instability which appears in systems with a van der Waals-like equations of state. Another possibility is the formation of the Bose condensate of a scalar field, in analogy with the Lee-Wick state, but in dilute matter [158].
We assume that growing unstable bosonic modes may form the Bose condensate described by a scalar boson complex field $\phi$. Such a scalar field can be introduced by means of the Hubbard–Stratonovich transformation or by a formal replacement of the contact interaction with an exchange of an artificial heavy scalar boson [158]. The effective Lagrangian density of the condensate field, taken in the simplest form $\phi=\phi_{0}e^{-i\omega_{c}t+i\vec{k}_{0}\vec{r}}$, can be presented as
@@
It was evaluated in [41, 42] and [75] in case of one-component fermion system. Energetically preferable is the state with $\omega_{c}=0$ as argued in Ref. [158]. Then we can estimate for $k_{0}\ll p_{\rm F}$,
@@
$a^{2}N\widetilde{\omega}^{2}_{s}(k_{0})\phi_{0}+\Lambda(0,k_{0})\,|\phi_{0}|^{2}\phi_{0}=0\,$$\quad$
$\widetilde{\omega}^{2}_{s}(k_{0})=|f_{0}(k_{0})|^{-1}-\Re\Phi(0,k_{0})\,.$ (53)
@@
$|\phi_{0}|^{2}$ $=-\frac{N\widetilde{\omega}^{2}_{s}(k_{0})}{a^{2}\lambda}\,\Big(1+\frac{k_{0}^{2}}{2p_{\rm F}^{2}}\Big)^{-1}\,$,
$E_{\rm B}$ $=-\frac{N^{2}\widetilde{\omega}^{4}_{s}(k_{0})}{2\lambda}\Big(1+\frac{k_{0}^{2}}{2p_{\rm F}^{2}}\Big)^{-1}.$ (54)
@@
Reference [158] employed a simple phenomenological parametrization for the volume part of the energy per nucleon, $\mathcal{E}_{N\rm}$ as an expansion in powers of $p_{\rm F}/m_{N}$ proposed in Ref. [137],